OpenStudy (anonymous):

Solve for w P = 2L + 2w

OpenStudy (help!!!!):

Isolate the equation where w is by itself. Subtract -2L on both sides and divide by 2.

OpenStudy (anonymous):

P = 2L + 2W 2L + P = 2W (2L + P)/2 = W

OpenStudy (chealyn98):

P = 2W +2L step is keep 2L in one side of equation , and move every thing else to the other side. In moving any quantity from one side to other, change the sign. ( equivalent of adding or subtracting equal quantity from both sides of equation). 2L = P - 2W divide both sides by 2: L = p/2 - w Step is the same as solving X in equation of 8 = 2X + 16 2X = 8 -16 = 8 X =4